# An Empirical Study of Testing Practices in Open Source AI Agent Frameworks and Agentic Applications

### v2

@@ 4 Related Work
4.1 Testing Ecosystem
4.1.1 Testing in OSS Ecosystem
Software testing is a foundational software engineering practice aimed at verifying quality and finding defects. A classic definition by Myers emphasizes the mindset: “Testing is the process of executing a program with the intent of finding errors” (Myers, 2006). Researchers have extensively studied how developers write and maintain tests in traditional software projects (e.g., Java, Python, JavaScript ecosystems), yielding common findings on test design, structure, and effectiveness. Meszaros’s xUnit Test Patterns catalog introduced common patterns in test code (Meszaros, 2007).
A practical lens for examining these techniques is the AAA (Arrange, Act, Assert) paradigm, which structures tests into clear phases of setup, execution, and verification (Wei et al., 2022a). While empirical studies show inconsistent adoption of specific design-level patterns (e.g., only 24% of projects applied certain maintainability-focused patterns (Gonzalez et al., 2017)), the structural AAA pattern has become ubiquitous, with 77% of Java unit tests adhering to its format to enhance readability (Wei et al., 2025). This review adopts the AAA structure to survey established testing techniques and foreshadow their application and limitations in the context of AI agents.
For the Arrange phase, we have seen different testing patterns in the OSS domain in the literature. For example, the use of test doubles (e.g., Mocks, Stubs, Fakes) eliminates the need for complex setup procedures, e.g., practitioners no longer need to set up databases for testing or simulate network failures through physical interventions like unplugging network cables (Zhu et al., 2023). Similarly, parameterized testing enables practitioners to execute the same test logic with multiple inputs and has been proven to run test functions 30 times faster than the traditional approach (Kampmann and Zeller, 2019).
In the Assert phase, OSS testing relies on verifiable outcomes. Studies have cataloged 12 assertion patterns, with equality checks being common (39.3% in JavaScript projects) (Zamprogno et al., 2022), mock assertion techniques, where practitioners often assert that certain calls were made on the mock (interaction-based verification), rather than asserting a returned value has also gained popularity with the test doubles is found in 41% of the test functions and 50% of the defects detected by mock assertion are missed by other assertion techniques (Zhu et al., 2025). The importance of assertion is also underscored by recent work on automatically generating asserts using ML and FMs (Fontes and Gay, 2023; Wang et al., 2024a). However, until now, it is not known how these structural patterns used for Arrange and verification patterns used for Assert are helping practitioners to test the agent frameworks and agentic applications.
4.1.2 Testing in ML Ecosystem
@@
4.2 Agent Ecosystem
4.2.1 Classical to FM-Based Agents
Building on the foundational BDI agent paradigm, which conceptualized agents as autonomous entities equipped with Beliefs, Desires, Intentions, and a plan library to support real-time decision-making, researchers expanded these concepts through formal models like OBDI that ground these mental constructs in computational semantics (Rao et al., 1995). More recent JaCaMo framework implements the core components of a classical BDI agent demonstrating a plan library (a set of pre-defined procedures for achieving goals), a belief base (a data structure to store beliefs), and an interpreter that runs a continuous reasoning cycle: perceive events, update beliefs, deliberate on desires to form intentions, and execute a relevant plan (Boissier et al., 2020). However, the advent of Foundation Models (FMs) has introduced a paradigm shift, replacing explicitly programmed logic of plan library with an FM as the core reasoning engine (Guo et al., 2024) The architecture of these FM-based agents typically integrates core components like an FM-based reasoning engine, clearly defined objectives, memory systems for historical context, an action module to interact with external tools and environments, and a reflective “Rethink Module” for continuous improvement enabling agents to utilize diverse tools such as APIs, robotic control systems, and code interpreters, significantly broadening their applicability across various domains (Cheng et al., 2024; Wang et al., 2024b; Masterman et al., 2024)
4.2.2 Agent Frameworks
To manage the complexity of building sophisticated FM-based agents, a new ecosystem of agent frameworks has emerged. For example, CAMEL framework leverages a role-playing paradigm with inception prompting, a prompting technique that enables agents to prompt each other to solve tasks, to generate large-scale conversational datasets for studying agent cooperation and “cognitive” behaviors (Li et al., 2023). On the other hand, AutoGen introduces conversable agents and conversation programming to abstract complex workflows into agent-to-agent dialogues, supporting mixed LLM, human, and tool interactions across diverse domains (Wu et al., 2023). MetaGPT encodes Standardized Operating Procedures (SOPs) into prompt sequences and uses an assembly-line model to decompose tasks among specialized agents, improving consistency and reducing hallucination cascades (Hong et al., 2023). Despite their strengths, their published evaluations focus on end-to-end application performance and benchmarking, leaving how they internally test core components (e.g., memory management, planning, tool invocation, and self-reflection) largely unexplored.
4.2.3 Evaluating Agent Frameworks and Agentic Systems
@@ 5.1.2 Metadata Filtering
To focus on well-maintained frameworks, we first filter repositories based on contributor count and popularity metrics. Prior research has highlighted that the number of contributors is a strong indicator of repository sustainability and adoption (Han et al., 2019). Therefore, we only retain repositories with at least two contributors.
Additionally, GitHub stars are commonly used as a proxy for repository popularity in empirical studies (Gonzalez et al., 2020; Munaiah et al., 2017; Openja et al., 2024) and we include only repositories with at least 1,000 stars. This step reduces the dataset to 67 repositories.
@@
As the primary objective of this study is to analyze testing practices, it is essential to ensure that all selected repositories contain at least one test file. In our study, we consider a file to be a test file if its filename started with the test_ prefix, following conventions from previous studies (Openja et al., 2024). We apply this filename-based test file detection technique and remove 5 repositories that do not contain any test files. This step results in 41 repositories.
@@ 5.1.4 Mining Test Functions
Table 2: Overview of AI agent frameworks and their GitHub statistics
@@
FoundationAgents/MetaGPT 56.4 146 240 964
microsoft/autogen 45.9 551 80 1041
crewAIInc/crewAI 32.9 248 50 588
agno-agi/agno 28.2 227 158 1697
huggingface/smolagents 20.1 157 18 463
openai/swarm 19.9 13 2 11
letta-ai/letta 16.8 135 30 650
eosphoros-ai/DB-GPT 16.8 146 73 728
TransformerOptimus/SuperAGI 16.4 73 114 466
raga-ai-hub/RagaAI-Catalyst 16.2 26 17 87
langchain-ai/langgraph 14.3 229 43 906
camel-ai/camel 12.9 148 221 1507
openai/openai-agents-python 11.4 96 61 563
pydantic/pydantic-ai 10.2 161 74 1096
QwenLM/Qwen-Agent 9.7 28 22 48
Upsonic/Upsonic 7.5 24 5 24
crestalnetwork/intentkit 6.4 15 2 17
livekit/agents 6.3 150 15 137
lavague-ai/LaVague 6.1 25 2 1
awslabs/agent-squad 6.0 23 22 312
superduper-io/superduper 5.1 48 65 318
kyegomez/swarms 4.9 44 54 645
MervinPraison/PraisonAI 4.8 20 33 241
AgentOps-AI/agentops 4.5 42 28 337
VRSEN/agency-swarm 3.7 19 5 78
SylphAI-Inc/AdalFlow 3.3 26 37 299
cheshire-cat-ai/core 2.8 86 45 197
Intelligent-Internet/$i$-agent 2.4 6 5 67
griptape-ai/griptape 2.3 37 256 1242
dot-agent/nextpy 2.3 12 88 828
InternLM/lagent 2.1 33 5 8
trypromptly/LLMStack 2.0 8 15 57
melih-unsal/DemoGPT 1.8 4 2 5
openlit/openlit 1.6 36 20 43
agentuniverse-ai/agentUniverse 1.5 27 30 70
fetchai/uAgents 1.4 60 15 136
patched-codes/patchwork 1.4 15 18 65
Div99/agent-protocol 1.2 15 11 22
@@
5.1.5 Stratified Random Sampling
To determine an appropriate sample size, we followed statistical practices from prior empirical software engineering studies (Openja et al., 2024). Using a 95% confidence level and a 5% margin of error, we calculate the required sample size using standard formulas for estimating proportions in finite populations, yielding a target of 377 test functions.
@@ 5.2.2 Filtering Toy Repositories
Table 4 presents the repository count at each filtering step. From our initial 7,397 repositories, these filters remove 6,958 repositories, and we prepare the remaining list of 439 repositories for the next step.
@@ 5.2.3 Mining and Deduplicating Test Functions
This de-duplication procedure removes the redundant test functions, resulting in the final dataset of 110,891 unique test functions sourced from agentic applications, all stored in our Elasticsearch index for further analysis.
@@ 5.2.4 Stratified Random Sampling
To mitigate this imbalance, we again apply stratified random sampling, treating each agentic application repository as a separate stratum and allocating sample size proportionally based on the number of test functions in each repository. Following the same statistical procedures used for the framework dataset, we target a 95% confidence level and a 5% margin of error. This results in a required sample size of 382 test functions, calculated using standard formulas for estimating proportions in finite populations. This approach ensures representative coverage of testing practices across agentic applications, while maintaining consistency with the sampling strategy used for framework repositories.
@@ 5.3 Card Sorting
We first apply hybrid card-sorting to identify emergent testing patterns, i.e., recurring structures, strategies, or practices embedded in the test code. This inductive approach allows patterns to emerge organically from the data while also allowing a set of predefined categories. (Zampetti et al., 2022). Next, to identify which parts of the agent architecture are being tested, we extract the subject under test (SUT) and apply a closed card sorting Zhao et al. (2024) to map those SUTs to the core agentic components defined in Section 3.
@@ 5.3.1 Hybrid Card Sorting for Testing Patterns
To analyze testing patterns in test functions extracted from AI agent frameworks and agentic applications, we apply hybrid card-sorting, following methodologies established in prior empirical studies (Zampetti et al., 2022). In a hybrid card-sorting, elements of both open and closed card sorting are combined, e.g., participants are given predefined categories into which to sort content (as in closed sorting), but are also free to create new categories when existing ones are insufficient (as in open sorting). Before detailing our card-sorting procedure, we first define what constitutes a testing pattern in this context and describe the underlying structure of unit tests.
@@
Testing patterns refer to these recurring strategies for structuring setup and verification logic in test functions. They improve the maintainability, diagnosability, and reliability of tests (Meszaros et al., 2003; Gonzalez et al., 2017). Practitioners may use mocking, patching, parameterization, or custom assertions to isolate behavior, simulate dependencies, or evaluate correctness.
@@
async def test_embedding(mocker): config = Config.default() mock_post = mocker.patch(’aiohttp.ClientSession.post’) mock_response = mocker.AsyncMock() mock_response.status = 200 data = await aread(Path(__file__).parent / ’../../data/openai/embedding.json’) mock_response.json.return_value = json.loads(data) mock_post.return_value.__aenter__.return_value = mock_response type(config.get_openai_llm()).proxy = mocker.PropertyMock(return_value=’http://mock.proxy’) llm_config = config.get_openai_llm() assert llm_config assert llm_config.proxy result = await oas3_openai_text_to_embedding(’Panda emoji’, openai_api_key=llm_config.api_key, proxy=llm_config.proxy) assert result assert result.model assert len(result.data) > 0 assert len(result.data[0].embedding) > 0 ArrangeActAssert
Figure 2: A Sample Test Function marked with Arrange, Act, Assert blocks.
@@
Because a single test function may exhibit multiple patterns, multi-label annotations were adopted. To ensure contextual understanding, raters consulted test code, SUT code, and README files. They also cross-referenced candidate patterns with those discussed in the five pre-reviewed studies when evaluating similarity or novelty.
@@
To address these issues, the raters developed a consensus taxonomy by merging redundancies, refining naming conventions in alignment with existing literature, and agreeing on consistent granularity levels. Once the taxonomy was finalized, the raters re-sorted the 15% dataset. Given the multi-label nature of this task, we measured inter-rater agreement using Jaccard similarity, following prior work on similar classification efforts (Parker et al., 2024). The agreement scores were 0.88 for agent framework test patterns and 0.85 for agentic application patterns, indicating high reliability of the coding scheme. These scores exceed the commonly accepted threshold of 0.80 for strong agreement, suggesting that the labels are well-defined and consistently interpretable. Based on this high level of agreement, the remainder of the dataset was coded by the first author using the finalized taxonomy.

### v3

@@ 4 Related Work
4.1 Testing in OSS Ecosystem
Software testing is a foundational software engineering practice aimed at verifying quality and finding defects (Myers, 2006). Researchers have extensively studied how developers write and maintain tests in traditional software projects (e.g., Java, Python, JavaScript ecosystems), yielding common empirical findings on test design patterns, structure, and effectiveness (Meszaros, 2007).
One widely observed technique is the AAA (Arrange-Act-Assert) pattern, which appears in 77% of Java unit tests to enhance readability (Wei et al., 2025). Within the arrange phase of this pattern, developers adopt recurring testing patterns in the OSS domain in the literature. For example, the use of test doubles (e.g., Mocks, Stubs, Fakes) enables isolation of the Subject Under Test from external dependencies, e.g., databases and remote services. For instance, prior work demonstrates how a data-access dependency can be stubbed to raise an exception on the first invocation and return a valid domain object on subsequent calls, allowing the login and recovery logic of a service to be verified without creating real database records (Zhu et al., 2023). Similarly, parameterized testing allows practitioners to execute the same test logic with multiple inputs and has been proven to run test functions 30 times faster than the traditional approach (Kampmann and Zeller, 2019).
In the assertion or verification phase, studies have cataloged 12 assertion patterns, with equality checks being common (39.3% in JavaScript projects) (Zamprogno et al., 2022). Mock assertion techniques have also gained popularity alongside the use of test doubles. In these tests, practitioners assert whether specific calls were made on a mock rather than asserting a returned value. Such assertions appear in 41% of test functions, and prior work shows that 50% of the defects detected by mock assertions are missed by other assertion techniques (Zhu et al., 2025). Recent works have further explored the use of ML and FMs to automatically generate assertions (Fontes and Gay, 2023; Wang et al., 2024). Despite the extensive work of testing patterns, it remains unclear how these well-established testing structures and patterns translate to agent frameworks and agentic applications that integrate non-deterministic model calls with procedural code.
4.2 Testing in ML Ecosystem
@@
4.3 Evaluating Agent Frameworks and Agentic Systems
@@ 5.1.2 Metadata Filtering
To focus on well-maintained frameworks, we first filter repositories based on contributor count and popularity metrics. Prior research has highlighted that the number of contributors is a strong indicator of repository sustainability and adoption (Han et al., 2019). Therefore, we only retain repositories with at least two contributors. While prior work often adopts higher thresholds (e.g., three contributors (Zou et al., 2019)) to remove small-scale projects, we intentionally relax this criterion to reflect the early stage of the agent ecosystem, which has only emerged in recent years and where many widely used frameworks are still being developed by small teams.
Additionally, GitHub stars are commonly used as a proxy for repository popularity in empirical studies (Gonzalez et al., 2020; Munaiah et al., 2017; Openja et al., 2024). We have observed the median star count of our dataset is 2,015 and choose the repositories with at least 1,000 stars following similar prior studies (Jebnoun et al., 2020) to preserve coverage of widely used and practically relevant agent frameworks. This step reduces the dataset to 67 repositories.
@@
As the primary objective of this study is to analyze testing practices, it is essential to ensure that all selected repositories contain at least one test file. In our study, we consider a file to be a test file if its filename starts with the test_ prefix, following conventions from previous studies (Jebnoun et al., 2020; Openja et al., 2024). Repositories without files following standard test naming conventions (e.g., without starting with test_ prefix) may require function-level analysis to identify test code, potentially complicating test extraction and analysis. Hence, we apply this prefix-based test file detection technique and remove 5 repositories that do not contain any test files. This step results in 41 repositories.
@@ 5.1.4 Mining Test Functions
  • $\bullet$
    The function is defined in a Python file intended for testing, as indicated by the filename pattern test_*.py.
  • $\bullet$
    The function is explicitly marked as a test through its name, which starts with the prefix test_ and is discoverable by common Python testing frameworks. [^2] https://docs.pytest.org/en/stable/how-to/unittest.html
@@
5.1.5 Stratified Random Sampling
To determine the appropriate sample size, we followed the statistical practices used in prior empirical software engineering studies and calculated the required sample size using Cochran’s formula with finite population correction (Chaokromthong et al., 2021), which is defined as:
$n=\frac{\dfrac{Z^{2}p(1-p)}{\varepsilon^{2}}}{1+\left(\dfrac{Z^{2}p(1-p)}{\varepsilon^{2}N}\right)}$
where $N$ is the population size, $Z$ is the $z$-score for the desired confidence level, $p$ is the estimated population proportion, and $\varepsilon$ is the margin of error. Following prior empirical studies (Openja et al., 2024), we set $Z=1.96$ for a 95% confidence level, $p=0.5$ as a conservative estimate, and $\varepsilon=0.05$. Substituting N=15,964N=15{,}964 yields an adjusted sample size of 377 (rounded up).
Table 2: Overview of AI agent frameworks and their GitHub statistics, including thousand lines of code (KLoC) as a proxy for repository size and complexity following prior studies (Barb et al., 2014).
@@ 5.1.5 Stratified Random Sampling
KLoC
FoundationAgents/MetaGPT 56.4 146 240 964 53.9
microsoft/autogen 45.9 551 80 1041 59.7
crewAIInc/crewAI 32.9 248 50 588 43.5
agno-agi/agno 28.2 227 158 1697 142.5
huggingface/smolagents 20.1 157 18 463 25.1
openai/swarm 19.9 13 2 11 1.2
letta-ai/letta 16.8 135 30 650 68.0
eosphoros-ai/DB-GPT 16.8 146 73 728 115.3
TransformerOptimus/SuperAGI 16.4 73 114 466 21.6
raga-ai-hub/RagaAI-Catalyst 16.2 26 17 87 6.4
langchain-ai/langgraph 14.3 229 43 906 68.1
camel-ai/camel 12.9 148 221 1507 92.8
openai/openai-agents-python 11.4 96 61 563 22.2
pydantic/pydantic-ai 10.2 161 74 1096 43.3
QwenLM/Qwen-Agent 9.7 28 22 48 12.7
Upsonic/Upsonic 7.5 24 5 24 3.2
crestalnetwork/intentkit 6.4 15 2 17 31.9
livekit/agents 6.3 150 15 137 40.1
lavague-ai/LaVague 6.1 25 2 1 7.3
awslabs/agent-squad 6.0 23 22 312 19.5
superduper-io/superduper 5.1 48 65 318 18.9
kyegomez/swarms 4.9 44 54 645 40.2
MervinPraison/PraisonAI 4.8 20 33 241 9.0
AgentOps-AI/agentops 4.5 42 28 337 22.8
VRSEN/agency-swarm 3.7 19 5 78 9.5
SylphAI-Inc/AdalFlow 3.3 26 37 299 22.5
cheshire-cat-ai/core 2.8 86 45 197 28.4
Intelligent-Internet/$i$-agent 2.4 6 5 67 6.7
griptape-ai/griptape 2.3 37 256 1242 44.7
dot-agent/nextpy 2.3 12 88 828 87.1
InternLM/lagent 2.1 33 5 8 7.6
trypromptly/LLMStack 2.0 8 15 57 5.5
melih-unsal/DemoGPT 1.8 4 2 5 4.3
openlit/openlit 1.6 36 20 43 7.1
agentuniverse-ai/agentUniverse 1.5 27 30 70 28.9
fetchai/uAgents 1.4 60 15 136 11.6
patched-codes/patchwork 1.4 15 18 65 11.5
Div99/agent-protocol 1.2 15 11 22 6.1
@@ 5.2.2 Filtering Toy Repositories
Table 4 presents the repository count at each filtering step. From our initial 7,397 repositories, these filters remove 6,958 repositories, and we prepare the remaining list of 439 repositories for the next step. To assess whether these repositories are representative of real-world agentic systems in terms of size and complexity, we measure their KLoC (thousands of lines of code), following prior empirical studies (Barb et al., 2014). The median repository size is 30 KLoC, suggesting that the retained projects are non-trivial and comparable to production-grade software systems.
To further examine the diversity of application domains covered by our dataset, we randomly sample 50 agentic applications from the final set and manually inspect their GitHub tags and descriptions. This analysis reveals a broad range of application categories, including generic assistants, software development and maintenance tools, medical assistants, crypto-related agents, and project management systems. Figure 2 summarizes this distribution, indicating that the dataset spans diverse domains rather than being concentrated in a single application area.
Figure 2: Distribution of agent application categories in our dataset. Each slice shows the proportion of agents assigned to a high-level category. Categories with more than one instance are reported individually: Generic Assistant (application, workflow manager or builder), Software Maintenance (monitoring, cloudops, review, testing, inferencing), Software Development (coding, documentation), Medicine, Crypto, and Project Management. All single-instance categories are aggregated into Others comprising assistants capable of data analysis, memory, blogger, ecommerce, hrmis, physics, and research.
@@ 5.2.3 Mining and Deduplicating Test Functions
This de-duplication procedure removes the redundant test functions, resulting in the final dataset of 110,891 unique test functions sourced from agentic applications, all stored in our Elasticsearch index for further analysis. In Table 5 we provide a count-level statistics for each of these steps.
Table 5: File- and function-level statistics of the AI agentic applications
Metric Count
# of total agentic application repositories 439
# of total files across all repositories 767,490
# of total test files across all repositories 40,831
# of total functions in all test files 789,343
# of extracted functions after de-duplication 110,891
@@ 5.2.4 Stratified Random Sampling
To mitigate this imbalance, we again apply stratified random sampling, treating each agentic application repository as a separate stratum and allocating sample size proportionally based on the number of test functions in each repository. Following the same Cochran formula (Chaokromthong et al., 2021; Ajibode et al., 2025) with finite population correction, and the same parameter settings ($Z=1.96$, $p=0.5$, and $\varepsilon=0.05$), we compute the initial required sample size for N=110,891N=110{,}891 as approximately 383. We then randomly select test functions without replacement within each repository stratum. During manual inspection, we found that one sampled test function had an empty implementation and therefore excluded it from further analysis, resulting in a final valid sample of 382 test functions. This approach ensures representative coverage of testing practices across agentic applications, while maintaining consistency with the sampling strategy used for framework repositories.
@@ 5.3 Card Sorting
We first apply hybrid card-sorting to identify emergent testing patterns, i.e., recurring structures, strategies, or practices embedded in the test code. This inductive approach allows patterns to emerge organically from the data while also allowing a set of predefined categories. (Zampetti et al., 2022). Next, to identify which parts of the agent architecture are being tested, we extract the subject under test (SUT) and apply a closed card sorting (Zhao et al., 2024) to map those SUTs to the core agentic components defined in Section 3.
@@ 5.3.1 Hybrid Card Sorting for Testing Patterns
To analyze testing patterns in test functions extracted from AI agent frameworks and agentic applications, we apply hybrid card-sorting, following methodologies established in prior empirical studies (Zampetti et al., 2022; Nayebi et al., 2018; Conrad and Tucker, 2019). In a hybrid card-sorting, elements of both open and closed card sorting are combined, e.g., participants are given predefined categories into which to sort content (as in closed sorting), but are also free to create new categories when existing ones are insufficient (as in open sorting). Before detailing our card-sorting procedure, we first define what constitutes a testing pattern in this context and describe the underlying structure of unit tests.
@@
1async def test_embedding(mocker): 2 config = Config.default() 3 mock_post = mocker.patch(’aiohttp.ClientSession.post’) 4 mock_response = mocker.AsyncMock() 5 mock_response.status = 200 6 data = await aread(Path(__file__).parent / ’../../data/openai/embedding.json’) 7 mock_response.json.return_value = json.loads(data) 8 mock_post.return_value.__aenter__.return_value = mock_response 9 type(config.get_openai_llm()).proxy = mocker.PropertyMock( 10 return_value=’http://mock.proxy’ 11 ) 12 llm_config = config.get_openai_llm() 13 assert llm_config 14 assert llm_config.proxy 15 result = await oas3_openai_text_to_embedding( 16 ’Panda emoji’, 17 openai_api_key=llm_config.api_key, 18 proxy=llm_config.proxy 19 ) 20 assert result 21 assert result.model 22 assert len(result.data) > 0 23 assert len(result.data[0].embedding) > 0 ArrangeActAssert
Figure 3: A sample test function marked with Arrange–Act–Assert blocks.
Testing patterns are these recurring strategies for structuring setup and verification logic in test functions. They improve the maintainability, diagnosability, and reliability of tests (Meszaros et al., 2003; Gonzalez et al., 2017). In practice, these strategies manifest as different techniques practitioners use; for example, they may use mocking, patching, parameterization, or custom assertions to isolate behavior, simulate dependencies, or evaluate correctness. To identify these patterns, we focus on analyzing the source code of the test functions and ground our findings through qualitative analysis, e.g., hybrid card sorting.
@@
A single test function can exhibit multiple patterns simultaneously. Hence, we employed a multi-label annotation strategy to account for the presence of such concurrent patterns within individual test functions. To guarantee a comprehensive understanding of the context, raters examined the test code, the System Under Test (SUT), and relevant README documentation. Additionally, they assessed the novelty of these patterns by cross-referencing them with the taxonomies established in the five selected studies.
@@
To address these issues, the raters developed a consensus taxonomy by merging redundancies, refining naming conventions in alignment with existing literature, and agreeing on consistent granularity levels. Once the taxonomy was finalized, the raters re-sorted the 15% dataset. Traditional agreement coefficients such as Fleiss’ Kappa and Krippendorff’s $\alpha$ assume that each item receives a single categorical label, whereas our annotation allows multiple labels per test function. Hence, applying these coefficients is not suitable for such multi-label cases (Marchal et al., 2022). Consequently, following prior multi-label annotation studies (Parker et al., 2024; Bitew et al., 2023), we use Jaccard similarity, which directly measures overlap between the two raters’ label sets. Specifically, each test function may be assigned multiple pattern labels by each rater; therefore, we treat annotations as sets of labels and compute the Jaccard coefficient as the ratio of the intersection of the two raters’ label sets to their union. To further explain, let $A_{i}$ and $B_{i}$ denote the sets of labels assigned by reviewer 1 and reviewer 2 to item $i$. The Jaccard similarity for item $i$ is defined as:
$J(A_{i},B_{i})=\frac{|A_{i}\cap B_{i}|}{|A_{i}\cup B_{i}|}$
where $|A_{i}\cap B_{i}|$ represents the number of labels assigned by both reviewers and $|A_{i}\cup B_{i}|$ represents the total number of unique labels assigned by either reviewer. The overall inter-rater agreement is computed as the average Jaccard similarity across all annotated items:
$J_{avg}=\frac{1}{N}\sum_{i=1}^{N}J(A_{i},B_{i})$
where $N$ is the number of items with at least one assigned label.
We observed agreement scores of 0.88 for agent framework patterns and 0.85 for agentic applications, both exceeding the 0.80 threshold for strong agreement. Though the original version of Krippendorff’s $\alpha$ does not handle multiple labels, a variant of Krippendorff’s $\alpha$ using the MASI distance (Passonneau, 2006) can be used to handle them. A prior study of disagreement in natural language inference used this MASI distance to calculate the inter-rater agreement (Jiang and Marneffe, 2022). Hence, to complement the Jaccard similarity, we also measured Krippendorff’s $\alpha$, with the MASI distance, which is 0.80 for agent frameworks and 0.75 for agentic applications. While Jaccard similarity captures raw set overlap between raters, Krippendorff’s $\alpha$ with MASI distance adjusts for chance agreement and partial label matches; thus, both metrics together provide a more comprehensive assessment of inter-rater agreement.